# TV homogenization inequalities

### v1

@@ 2 Proofs
2.1 Proof of Theorem 1
We will prove a sharper, symmetric bound, which immediately implies the one in Theorem 1:
Theorem 6.
@@ 2.1 Proof of Theorem 1
by a simple telescoping argument.
Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman). If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
@@
Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Endrei-Schoenberg-Whitney theorem [15, Section 4] that the Poisson binomial has a log-concave, and hence unimodal, law.
To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
@@
Then $T_{i}(t)$ is independent of $X_{i}(t)$ and $S(t)=T_{i}(t)+X_{i}(t)$.
Conditioning on $X_{i}(t)$ gives
@@
Indeed, if $x\leq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{\mathrm{BCV}}/\sqrt{x+1}\geq C_{\mathrm{BCV}}/\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}=1$. If $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, then the left-hand side equals $\eta_{\mathrm{BCV}}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$.
Applying (8) with $x=\operatorname{Var}(S(t))$ gives
@@
Taking $\sup_{A}$ proves (6). ∎
2.2 Proof of Theorem 2
We begin with an abstract result showing that if a positive measure $g$ on $\mathbb{Z}$ has total mass $G$ and its variance-like spread about $\mu$ is controlled by $J<\infty$, this implies a lower bound on the mass of the heaviest atom.
Lemma 7 (Pigeonhole).
@@ 2.2 Proof of Theorem 2
$\sup_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{3\,G^{3/2}}{16\sqrt{J}+4\sqrt{G}}.$
Proof.
Put $R:=2\sqrt{{J}/{G}}$ and define
$A:=\{k\in\mathbb{Z}:\ |k-\mu|\leq R\},\qquad A^{\mathrm{c}}:=\mathbb{Z}\setminus A.$
A Chebyshev-type argument shows that at least $\tfrac{3}{4}$ of the mass of $g$ lies in $A$. Indeed, on $A^{\mathrm{c}}$ we have $(k-\mu)^{2}>R^{2}$, so
$J=\sum_{k}(k-\mu)^{2}g(k)\geq\sum_{k\in A^{\mathrm{c}}}(k-\mu)^{2}g(k)>R^{2}\sum_{k\in A^{\mathrm{c}}}g(k),$
hence
$\sum_{k\in A^{\mathrm{c}}}g(k)\leq\frac{J}{R^{2}}=\frac{J}{4J/G}=\frac{G}{4}.$
Therefore
$\sum_{k\in A}g(k)=G-\sum_{k\in A^{\mathrm{c}}}g(k)\geq\frac{3G}{4}.$
Next, we extract the “heavy” atom in $A$ via a cardinality bound and the pigeonhole principle:
$|A|\leq 2R+1=4\sqrt{\frac{J}{G}}+1,$
and so
$\sup_{k\in\mathbb{Z}}g(k)\ \geq\ \sup_{k\in A}g(k)\ \geq\ \frac{1}{|A|}\sum_{k\in A}g(k)\ \geq\ \frac{(3G/4)}{4\sqrt{J/G}+1}.$
The latter simplifies to our claimed bound. ∎
@@
$G:=\Delta=m_{\textup{{\bf p}}}-m_{\textup{{\bf q}}}=\mathbb{E}X-\mathbb{E}Y,$
where $X:=S_{\textup{{\bf p}}},Y:=S_{\textup{{\bf q}}}$ and $m_{\textup{{\bf p}}}:=\mathbb{E}X,m_{\textup{{\bf q}}}:=\mathbb{E}Y$. Moreover, for any $k$,
@@
Finally, $J=\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}g(k)$ is defined as in Lemma 7. The technical core of the argument hinges on upper bounding $J$:
Lemma 8.
Let $X,Y,\sigma_{\textup{{\bf p}}}^{2}=\operatorname{Var}(X),g,m_{\textup{{\bf p}}},\Delta,J$ be defined as above. Then
@@
In this formulation, the functional $J=\sum_{k}(k-m_{\textup{{\bf p}}})^{2}g(k)$ becomes
$J$ $=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t.$
@@
$\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)=\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr],$
@@
$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
We bound $\mathbb{E}_{t}[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}]$ uniformly in $t,i$ via the decomposition $\mathbb{E}[(Z-\alpha)]^{2}=\operatorname{Var}(Z)+(\mathbb{E}Z-\alpha)^{2}$, where $Z=T_{i}(t)+1$ and $\alpha=m_{\textup{{\bf p}}}$.
To bound the variance term, note that f(x)=x(1x)f(x)=x(1-x) is $1$-Lipschitz on $[0,1]$, and so
@@
$\mathbb{E}_{t}[S(t)]=m_{\textup{{\bf q}}}+t\Delta,$
and $m_{\textup{{\bf p}}}=m_{\textup{{\bf q}}}+\Delta$, so $\mathbb{E}_{t}[S(t)]-m_{\textup{{\bf p}}}=-(1-t)\Delta$. Also $r_{i}(t)=q_{i}+t\Delta_{i}$, so
$\mathbb{E}_{t}[T_{i}(t)]=m_{\textup{{\bf q}}}+t\Delta-(q_{i}+t\Delta_{i}),$
@@
$M_{i}(t):=\mathbb{E}_{t}[T_{i}(t)+1-m_{\textup{{\bf p}}}]=1-q_{i}-\Delta+t(\Delta-\Delta_{i}).$
@@
$\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr]$ $=$ $\operatorname{Var}_{t}(T_{i}(t))+M_{i}(t)^{2}$
@@
Corollary 9 (Proof of Theorem 2).
@@
for some universal constant $c\geq\frac{1}{12}$.
@@
$\Delta=\sum_{k}g(k),\qquad J=\sum_{k}(k-m_{\textup{{\bf p}}})^{2}g(k),\qquad m_{\textup{{\bf p}}}=\mathbb{E}S_{\textup{{\bf p}}}.$
By (11), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\geq\sup_{k\in\mathbb{Z}}g(k)$. By Lemma 8,
@@
Apply Lemma 7 with $G=\Delta$ to obtain
$\sup_{k}g(k)\;\geq\;\frac{3\Delta}{16\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})}+4}.$
We compare $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}}$ with $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$ and consider the two cases. If $\Delta\leq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\leq\sigma_{\textup{{\bf p}}}^{2}+1$, so
$\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2(\sigma_{\textup{{\bf p}}}^{2}+1),$
and
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{3\Delta}{16\sqrt{2}\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1)}+4}=\frac{3\Delta}{32\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+4}$
$\geq\frac{3\Delta}{36\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}=\frac{\Delta}{12\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}.$
If $\Delta\geq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\geq\sigma_{\textup{{\bf p}}}^{2}+1$, so
$\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2\Delta^{2},$
and
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{3\Delta}{16\sqrt{2}\sqrt{2\Delta^{2}}+4}=\frac{3\Delta}{32\Delta+4}\geq\frac{3\Delta}{36\Delta}=\frac{1}{12}.$
@@ 2.3 Proof of Theorem 3
and the second in the form $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$. Given these, the claim is an immediate consequence of Theorem 2.
2.4 Proof of Lemma 4
@@
$\Delta_{I}(m):=\operatorname{TV}(\operatorname{Bin}(m,\bar{p}_{I}),\operatorname{Bin}(m,\bar{q}_{I})),\qquad\Delta_{J}(k):=\operatorname{TV}(\operatorname{Bin}(k,\bar{p}_{J}),\operatorname{Bin}(k,\bar{q}_{J})).$
@@
$\delta_{N}\leq\mathbb{E}[\Delta_{I}(M)]+\mathbb{E}[\Delta_{J}(n-M)].$
@@
Apply this with $f=\Delta_{I}$ and $m_{0}=|I|$ (note $\mathbb{E}M=nw=|I|$) to get
$\mathbb{E}[\Delta_{I}(M)]\leq 2\Delta_{I}(|I|)=2\delta_{I},$
and with $f=\Delta_{J}$ and $m_{0}=|J|$ (since $\mathbb{E}(n-M)=|J|$) to get
$\mathbb{E}[\Delta_{J}(n-M)]\leq 2\Delta_{J}(|J|)=2\delta_{J}.$
@@
Acknowledgments.
Some of the proofs were assisted by ChatGPT Pro; we thank Sébastien Bubeck for making this version available to us. We also thank Mark Sellke and Nikita Zhivotovskiy for pointing out errors in earlier versions.

### v2

@@ 2 Proofs
2.1 Proof of Theorem 1.1
We will prove a more general, symmetric bound, which immediately implies the one in Theorem 1.1. The constant $2C_{\mathrm{BCV}}$ in the latter’s bound could be sharpened by a factor of $2$ via a direct (asymmetric) proof. L. Mattner [12] has brought to our attention the result of [21], which upper bounds $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ by
$\frac{5|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{\sqrt{\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}+\frac{12|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{{1+\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}.$
This estimate has the attractive property of being topologically equivalent to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ in the sense that the two converge to 0 on precisely the same set of sequence pairs. It is currently an open question whether a version of our $\Phi$ functional could be made topologically equivalent to $\operatorname{TV}$ as well.
Theorem 2.1.
@@
by a simple telescoping argument. [^3] M. Lutz [12] informs us that this result is contained in [2] with a slightly worse constant. Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman). If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
@@
Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Edrei-Schoenberg-Whitney theorem [20, Section 4] that the Poisson binomial
270]
has a log-concave, and hence unimodal, law. To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
@@
Then $T_{i}(t)$ is independent of $X_{i}(t)$ and $S(t)=T_{i}(t)+X_{i}(t)$. Conditioning on $X_{i}(t)$ gives
@@
Indeed, if $x\leq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{\mathrm{BCV}}/\sqrt{x+1}\geq C_{\mathrm{BCV}}/\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}=1$. If $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, then the left-hand side equals $\eta_{\mathrm{BCV}}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$. Applying (8) with $x=\operatorname{Var}(S(t))$ gives
@@
Taking $\sup_{A}$ proves (2.1). ∎
2.2 Proof of Theorem 1.2
We make use of a concentration-variance inequality [13, Disp.(18)], which immediately implies the following “pigeonhole” lemma: [^4] We thank M. Lutz [12] for bringing this to our attention and improving our original lower bound by a factor of $4/3$.
Lemma 2.2.
@@
$\sup_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{4\sqrt{J}+\sqrt{G}}.$
Proof (sketch).
Normalize $g$ to be a probability measure $P$ via $P(\left\{k\right\})=g(k)/G$ and let $X$ be distributed according to $P$. Using the variational characterization of variance, $Var(X)\leq J/G$.
By Mattner–Schulz [13, Disp. (18)], for any law with variance $\sigma^{2}$ and any $h>0$,
$\sup_{x\in\mathbb{R}}P\bigl((x,x+h)\bigr)\;\geq\;\frac{h}{\sqrt{h^{2}+12\sigma^{2}}}.$
Apply this with $h=1$ and $\sigma^{2}=\operatorname{Var}(X)$. Since $P$ is supported on $\mathbb{Z}$, any interval of length $1$ contains at most one atom, and hence
$\sup_{x\in\mathbb{R}}P((x,x+1))=\max_{k\in\mathbb{Z}}P(\{k\})=\frac{1}{G}\max_{k\in\mathbb{Z}}g(k).$
Therefore,
$\frac{1}{G}\max_{k}g(k)\;\geq\;\frac{1}{\sqrt{1+12\operatorname{Var}(X)}}.$
Using $\operatorname{Var}(X)\leq J/G$,
$\max_{k}g(k)\;\geq\;\frac{G}{\sqrt{1+12(J/G)}}=\frac{G^{3/2}}{\sqrt{G+12J}}.$
Finally, since $\sqrt{G+12J}\leq\sqrt{G}+4\sqrt{J}$, we conclude
$\max_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{\,4\sqrt{J}+\sqrt{G}\,}$,
which is the claimed bound. ∎
@@
$G:=\Delta=\mu_{\textup{{\bf p}}}-\mu_{\textup{{\bf q}}}=\mathbb{E}X-\mathbb{E}Y,$
where $X:=S_{\textup{{\bf p}}},Y:=S_{\textup{{\bf q}}}$ and $\mu_{\textup{{\bf p}}}:=\mathbb{E}X,\mu_{\textup{{\bf q}}}:=\mathbb{E}Y$. Moreover, for any $k$,
@@
Finally, $J=\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}g(k)$ is defined as in Lemma 2.2. The technical core of the argument hinges on upper bounding $J$:
Lemma 2.3.
Let $X,Y,\sigma_{\textup{{\bf p}}}^{2}=\operatorname{Var}(X),g,\mu_{\textup{{\bf p}}},\Delta,J$ be defined as above. Then
@@
In this formulation, the functional $J=\sum_{k}(k-\mu_{\textup{{\bf p}}})^{2}g(k)$ becomes
$J$ $=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t.$
@@
$\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)=\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr],$
@@
$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
We bound $\mathbb{E}_{t}[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}]$ uniformly in $t,i$ via the decomposition $\mathbb{E}[(Z-\alpha)]^{2}=\operatorname{Var}(Z)+(\mathbb{E}Z-\alpha)^{2}$, where $Z=T_{i}(t)+1$ and $\alpha=\mu_{\textup{{\bf p}}}$. To bound the variance term, note that f(x)=x(1x)f(x)=x(1-x) is $1$-Lipschitz on $[0,1]$, and so
@@
$\mathbb{E}_{t}[S(t)]=\mu_{\textup{{\bf q}}}+t\Delta,$
and $\mu_{\textup{{\bf p}}}=\mu_{\textup{{\bf q}}}+\Delta$, so $\mathbb{E}_{t}[S(t)]-\mu_{\textup{{\bf p}}}=-(1-t)\Delta$. Also $r_{i}(t)=q_{i}+t\Delta_{i}$, so
$\mathbb{E}_{t}[T_{i}(t)]=\mu_{\textup{{\bf q}}}+t\Delta-(q_{i}+t\Delta_{i}),$
@@
$M_{i}(t):=\mathbb{E}_{t}[T_{i}(t)+1-\mu_{\textup{{\bf p}}}]=1-q_{i}-\Delta+t(\Delta-\Delta_{i}).$
@@
$\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]$ $=$ $\operatorname{Var}_{t}(T_{i}(t))+M_{i}(t)^{2}$
@@
Corollary 2.4 (Proof of Theorem 1.2).
@@
for some universal constant $c\geq\frac{1}{9}$.
@@
$\Delta=\sum_{k}g(k),\qquad J=\sum_{k}(k-\mu_{\textup{{\bf p}}})^{2}g(k),\qquad\mu_{\textup{{\bf p}}}=\mathbb{E}S_{\textup{{\bf p}}}.$
By (11), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\geq\sup_{k\in\mathbb{Z}}g(k)$. By Lemma 2.3,
@@
Apply Lemma 2.2 with $G=\Delta$ to obtain
$\sup_{k}g(k)\;\geq\;\frac{\Delta}{4\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})}+1}.$
We compare $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}}$ with $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$ and consider the two cases. If $\Delta\leq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\leq\sigma_{\textup{{\bf p}}}^{2}+1$, so $\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2(\sigma_{\textup{{\bf p}}}^{2}+1),$ and
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{\Delta}{4\sqrt{2}\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1)}+1}=\frac{\Delta}{8\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+1}\geq\frac{\Delta}{9\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}.$
If $\Delta\geq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\geq\sigma_{\textup{{\bf p}}}^{2}+1$, so $\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2\Delta^{2},$ and
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{\Delta}{4\sqrt{2}\sqrt{2\Delta^{2}}+1}=\frac{\Delta}{8\Delta+1}\geq\frac{1}{9}.$
@@
and the second in the form $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$. Given these, the claim is an immediate consequence of Theorem 1.2.
2.4 Proof of Lemma 1.4
@@
$\tau_{I}(m):=\operatorname{TV}(\operatorname{Bin}(m,\bar{p}_{I}),\operatorname{Bin}(m,\bar{q}_{I})),\qquad\tau_{J}(k):=\operatorname{TV}(\operatorname{Bin}(k,\bar{p}_{J}),\operatorname{Bin}(k,\bar{q}_{J})).$
@@
$\delta_{N}\leq\mathbb{E}[\tau_{I}(M)]+\mathbb{E}[\tau_{J}(n-M)].$
@@
Apply this with $f=\tau_{I}$ and $m_{0}=|I|$ (note $\mathbb{E}M=nw=|I|$) to get
$\mathbb{E}[\tau_{I}(M)]\leq 2\tau_{I}(|I|)=2\delta_{I},$
and with $f=\tau_{J}$ and $m_{0}=|J|$ (since $\mathbb{E}(n-M)=|J|$) to get
$\mathbb{E}[\tau_{J}(n-M)]\leq 2\tau_{J}(|J|)=2\delta_{J}.$
@@
Acknowledgements
Some of the proofs were assisted by ChatGPT Pro. We thank Sébastien Bubeck, Mark Sellke and Nikita Zhivotovskiy for pointing out errors in earlier versions. We are grateful to Lutz Mattner and Bero Roos for detailed and insightful comments on the manuscript.
Appendix A Homogenization is a Markov Kernel in the symmetric case
We show that for symmetric $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ — i.e., those satisfying $\textup{{\bf p}}=\textup{{\bf 1}}-\textup{{\bf q}}$, homogenization can be realized by a Markov kernel. We first consider the case $n=2$:
Lemma A.1 (averaging centered Bernoullis).
Let $\alpha,\beta\in(0,1)$ and put $\gamma=(\alpha+\beta)/2$. Then there exists a randomized function $F_{\alpha,\beta}:\left\{0,1\right\}^{2}\to\left\{0,1\right\}^{2}$ with the following property:
$X\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(X)\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),$
$Y\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(Y)\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right).$
Proof.
As in the proof of [9, Lemma 2.3], it suffices to write down the stochastic vectors representing $X,Y$ and $\hat{X}\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),\hat{Y}\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)$, the stochastic matrix $M$ representing $F_{\alpha,\beta}$, and to verify that $\hat{X}=XM$ and $\hat{Y}=YM$. For readability, we mildly abuse notation and blur the distinction between the random variable $Z\sim\operatorname{Ber}(p)\otimes\operatorname{Ber}(q)$ and its stochastic vector:
$Z=\left(\mathbb{P}(Z=00),\mathbb{P}(Z=01),\mathbb{P}(Z=10),\mathbb{P}(Z=11)\right)$ $=$ $\left((1-p)(1-q),(1-p)q,p(1-q),pq\right).$
With this convention, the four stochastic vectors are:
$X$ $=$ $\left(\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right)\right),$
$Y$ $=$ $\left(\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right)\right),$
$\hat{X}$ $=$ $\left(\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right)\right),$
$\hat{Y}$ $=$ $\left(\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right)\right).$
With $\eta:=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}$, we define the $4\times 4$ row-stochastic matrix $M$:
$M=\left(\begin{array}[]{cccc}1&0&0&0\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ 0&0&0&1\end{array}\right).$ (19)
Verifying the identities $\hat{X}=XM$ and $\hat{Y}=YM$ is tedious but quite straightforward. Obviously, the rows of $M$ sum to $1$. It only remains to show that the entries of $M$ are nonnegative, or equivalently, that $\eta=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}\leq 1/2$. The latter is equivalent to ${\alpha^{2}-2\alpha\beta+\beta^{2}}\leq{4(1-\alpha\beta)}$. Expanding and collecting like terms yields the equivalent claim $4-(\alpha+\beta)^{2}\geq 0$, which obviously holds for $\alpha,\beta\in(0,1)$. ∎
For general $n$, the pairwise averaging kernel $M=M(\alpha_{i},\beta_{i})$ defined in (19) can be applied repeatedly to the pairs $(i,j)$. Although the kernel itself changes with each iteration, a standard compactness argument shows that for any symmetric pair $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ there exists a Markov kernel $M_{\textup{{\bf p}},\textup{{\bf q}}}$ simultaneously achieving $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Ber}(\bar{p})^{\otimes n}$ and $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf q}})\mapsto\operatorname{Ber}(\bar{q})^{\otimes n}$.